16(7n+8)=-8(5n-(65))
this seems difficult but i came up with
240n=2600n
can you explain what i did wrong

Answer by wuwei96815(245)

16(7n+8)=-8(5n-(65))
Remove inner parenthesis.
16(7n+8)=-8(5n-65)
Remove outer parentheses.
112n+128=-40n+520
Add 40n to both sides.
152n+128=520
Add -128 to both sides.
152n=392
Divide both sides by 152.
n=2.58

Is that the right answer?